Groupe de contact du Fonds National de la Recherche Scientifique.

University of Mons-Hainaut

May 21, 2002

Program

10h00 Welcome Coffee

10h15-11h30 E. Jaligot (C.N.R.S., Paris 7, F.)

“Groups of finite Morley rank (to be confirmed)”

11h30-12h45 E.I. Khukhro

(Cardiff University, U.K., Novosibirsk Institute of Mathematics, R.)

“Restricted Burnside type problems for groups with operators”

12h45-14h15 Lunch

14h15-15h30 J.S. Wilson (Birmingham University, U.K.)

“Invisible subgroups of finite index”

15h30-16h45 N.Y. Makarenko (Novosibirsk Institute of Mathematics, R.)

“Largely splitting automorphisms”

16h45-17h15 Coffee break

17h15-18h30 R. Grigorchuk (Steklov Institute, Moscow, R.)

“Groups generated by finite automata and amenability”

In the evening, a dinner in town will be organized.

The meeting will take place in room 0A07 of the department of mathematics of Mons-Hainaut University (“Le Pentagone”, 6, Avenue du Champ de Mars, 7000 Mons.)

If you are interested and want to receive further information, please contact Françoise Point (email: point@logique.jussieu.fr), or Christian Michaux (email: chrmich@sun1.umh.ac.be), or Khaled Jaber (email: jaber@umh.ac.be). Please register for the lunch and the dinner before May 14, 2002. On-line information: http://www.umh.ac.be/math/logic